OpenStudy (anonymous):

find the length of each segment ac and bd a=-8 b= -6 c= 1 D= 3

OpenStudy (anonymous):

AC = -8(1) AC = -8 BD = -6(3) BD = -18

OpenStudy (anonymous):

To find the length of a segment, we take the absolute value of the difference of the segment's end-points on the number line. AC = |c-a| = |1-(-8)| = |1+8| = |9| = 9 BD= |d-b| = |3-(-6)| = |3+6| = |9| = 9